OpenStudy (anonymous):

Choose the equation of the line passing through the point (-3, -1) and perpendicular to y = x + 3. a. y = -2x - 7 b. y = -2x +5 c.y = 2x + 5 d. y = 2x - 7

OpenStudy (nubeer):

gradient of perpendicula = -1/m gradient = -1 now apply formula y-y1=m(x-x1)